# Weak Adversarial Neural Pushforward Method for Fractional Fokker-Planck Equations

### 13 Mar 2026 v1

@@ Abstract
We extend the Weak Adversarial Neural Pushforward Method (WANPM) to fractional Fokker-Planck equations (fFPE), in which the classical Laplacian diffusion operator is replaced by the fractional Laplacian $(-\Delta)^{\alpha/2}$ for $\alpha\in(0,2]$. The solution distribution is represented not as an explicit probability density function but as the pushforward of a simple base distribution through a time-parameterized neural network $F_{\bm{\vartheta}}(t,\mathbf{x}_{0},\mathbf{r})$, which enforces the initial condition exactly by construction. The weak formulation of the fFPE is discretized via Monte Carlo sampling entirely without temporal discretization, and the resulting min-max objective is optimized adversarially against a set of plane-wave test functions. A key computational advantage is that plane waves are eigenfunctions of the fractional Laplacian, so $(-\Delta_{x})^{\alpha/2}f=|\mathbf{w}|^{\alpha}f$ is computed exactly and at no additional cost for any $\alpha$. We validate the method on a one-dimensional fractional Fokker-Planck equation with a quadratic confining potential and $\alpha=1.5$, using a particle simulation based on symmetric $\alpha$-stable Lévy increments as a benchmark. The learned solution faithfully reproduces the transient probability distribution over $t\in[0,2]$, and robust statistics (median, IQR, MAD, and percentile bands) confirm close agreement with the particle simulation, while standard deviation comparisons highlight why second-moment metrics are inappropriate for heavy-tailed ($\alpha<2$) distributions. The present work lays the groundwork for mesh-free, scalable solvers for anomalous diffusion and fractional kinetics in high dimensions.
@@ 1 Introduction
Solving the fFPE in high dimensions poses formidable challenges. Classical finite-element or finite-difference discretizations of the fractional Laplacian require dense stiffness matrices whose assembly and storage become prohibitive even in modest dimensions. Spectral methods are effective in simple geometries but do not readily extend to curved or unbounded domains. Deep learning approaches, which have recently achieved remarkable success for classical high-dimensional PDEs, must contend with two additional obstacles in the fractional setting: (i) pointwise neural-network representations of the solution density require computing the fractional Laplacian of the network output, which involves non-local integral operators and is expensive to evaluate; and (ii) the solution distribution may have heavy algebraic tails that a neural network density must resolve over an exponentially large effective domain.
In a companion paper [3] we introduced the Weak Adversarial Neural Pushforward Method (WANPM) for classical Fokker-Planck equations. The core idea is to represent the solution distribution as the pushforward of a tractable base distribution through a neural map $F_{\bm{\vartheta}}$, and to train $F_{\bm{\vartheta}}$ adversarially against a family of plane-wave test functions so that the induced distribution satisfies the weak form of the FPE. This approach avoids explicit density evaluation, density normalization, and second-order backpropagation through the solution network.
@@
    We derive the weak formulation of the time-dependent fFPE and show how it can be evaluated as Monte Carlo expectations with respect to the pushforward distribution, without discretizing time.
@@
    We describe the pushforward architecture and the fractional Laplacian eigenfunction property that makes plane-wave test functions exact and efficient for the fractional operator.
@@
    We implement and validate the method on a one-dimensional quadratic confining potential with $\alpha=1.5$, comparing the learned transient distribution against a particle simulation using symmetric $\alpha$-stable Lévy increments.
@@
The remainder of the paper is organized as follows. Section 2 presents the fractional Fokker-Planck equation and its weak formulation. Section 3 describes the neural pushforward architecture and the adversarial training procedure. Section 4 presents the particle simulation used as a ground-truth benchmark. Section 5 reports the numerical experiments. Section 6 discusses limitations and future directions.
@@ 2.1 Strong Form
For the numerical experiments of this paper we take
$b(x)=-kx,\qquad k=1,$ (3)
corresponding to the gradient drift of a quadratic (harmonic) confining potential $V(x)=\tfrac{1}{2}kx^{2}$. With standard Brownian motion ($\alpha=2$), the solution is a Gaussian whose mean and variance can be tracked analytically. For $\alpha<2$ no closed-form expression for the density is available in general, and the solution develops algebraically decaying tails characteristic of $\alpha$-stable distributions.
Remark 1 (Comparison with the fractional diffusion normalization).
Different conventions for the prefactor of $(-\Delta)^{\alpha/2}$ appear in the literature. Our convention sets the coefficient to unity, so that the characteristic function of the Lévy increment over a unit time interval is $e^{-|\xi|^{\alpha}}$. The fractional diffusion coefficient can be absorbed into the wavenumber normalization of the test functions without loss of generality.
2.2 Weak Formulation
Let $f(t,x)$ be a smooth test function that decays sufficiently rapidly as $|x|\to\infty$. Multiplying (2) by $f$ and integrating over $\mathbb{R}\times[0,T]$ by parts in both space and time yields
@@ 2.2 Weak Formulation
The identity $(-\Delta_{x})^{\alpha/2}f=|\xi|^{\alpha}\hat{f}$ ensures that $\mathcal{L}f$ is well-defined for any smooth $f$, and the integration-by-parts step is valid whenever $\rho$ has sufficient integrability.
@@
Equation (4) decomposes into three Monte Carlo-evaluable expectations: a terminal expectation at $t=T$, an initial-condition expectation at $t=0$ (which can be computed from samples of $\rho_{0}$ independently of the learned network), and an interior expectation integrated uniformly over $[0,T]$. This structure is identical to that of the classical FPE [3] up to the replacement of $\partial^{2}/\partial x^{2}$ by $(-\Delta_{x})^{\alpha/2}$ in the operator $\mathcal{L}$.
@@ 2.3 Eigenfunction Property of Plane-Wave Test Functions
The central computational observation motivating our choice of test functions is the following. A spatial plane wave $\phi_{\mathbf{w}}(x)=\sin(\mathbf{w}\cdot x+\varphi)$ or $\cos(\mathbf{w}\cdot x+\varphi)$ satisfies
@@
As a consequence, the fractional part of $\mathcal{L}f^{(k)}$ costs nothing extra to evaluate compared with classical diffusion: it is simply a multiplication by the scalar $|\mathbf{w}^{(k)}|^{\alpha}$. This makes the fFPE loss computation as cheap as the classical FPE loss computation, regardless of $\alpha$.
@@ 3 Weak Adversarial Neural Pushforward Method
3.1 Pushforward Network Architecture
@@
$F_{\bm{\vartheta}}:[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{d}\to\mathbb{R},$ (7)
@@
where $\tilde{F}_{\bm{\vartheta}}:\mathbb{R}^{1+n+d}\to\mathbb{R}^{n}$ is a fully connected feedforward network with three hidden layers of width 128 and Tanh activation functions. At $t=0$, the map reduces to the identity on $x_{0}$, so $F_{\bm{\vartheta}}(0,x_{0},\mathbf{r})=x_{0}\sim\rho_{0}$ for any values of the network parameters. The $\sqrt{t}$ prefactor is appropriate for diffusion-driven processes: the typical displacement of a Lévy process over time $t$ scales as $t^{1/\alpha}$, and for $\alpha$ near 2 (and exactly for $\alpha=2$) this is well-approximated by $\sqrt{t}$ at short times.
The base distribution dimension $d=5$ was used in our experiments, providing the network with more latent degrees of freedom than the output dimension, which empirically improves the network’s ability to represent heavy-tailed distributions.
Remark 3 (Overparameterized base distributions).
Allowing the base dimension $d$ to exceed the target dimension $n$ precludes the use of invertible (normalizing flow) architectures, which require $d=n$. This is not a limitation but a feature: the extra latent dimensions provide additional representational capacity without requiring Jacobian-determinant computations.
@@ 3.2 Plane-Wave Test Functions
where (12) follows from the eigenfunction property (6). No automatic differentiation through the test function is required.
The operator evaluated on the test function is therefore
@@
3.3 Monte Carlo Discretization of the Weak Form
The three terms in the weak residual (4) are each approximated by Monte Carlo estimates.
Terminal term.
Sample $M_{T}$ tuples $(x_{0},\mathbf{r})$ with $x_{0}\sim\rho_{0}$ and $\mathbf{r}\sim\mathcal{N}(0,I_{d})$, and evaluate the pushforward at $t=T$:
$\hat{E}_{T}^{(k)}=\frac{1}{M_{T}}\sum_{m=1}^{M_{T}}f^{(k)}\!\left(T,\,F_{\bm{\vartheta}}(T,x_{0}^{(m)},\mathbf{r}^{(m)})\right).$ (14)
Initial condition term.
Sample $M_{0}$ particles $\tilde{x}^{(i)}\sim\rho_{0}=\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ and evaluate the test function at $t=0$:
$\hat{E}_{0}^{(k)}=\frac{1}{M_{0}}\sum_{i=1}^{M_{0}}f^{(k)}\!\left(0,\,\tilde{x}^{(i)}\right).$ (15)
Note that $f^{(k)}(0,x)=\sin(w^{(k)}x+b^{(k)})$, so this term is computed from the known initial distribution and does not involve the pushforward network.
Interior term.
Sample $M$ triples $(t^{(m)},x_{0}^{(m)},\mathbf{r}^{(m)})$ with $t^{(m)}\sim\mathcal{U}[\epsilon,T]$, $x_{0}^{(m)}\sim\rho_{0}$, and $\mathbf{r}^{(m)}\sim\mathcal{N}(0,I_{d})$, and form the pushforward $x^{(m)}=F_{\bm{\vartheta}}(t^{(m)},x_{0}^{(m)},\mathbf{r}^{(m)})$:
$\hat{E}^{(k)}=\frac{T}{M}\sum_{m=1}^{M}\left[\frac{\partial f^{(k)}}{\partial t}+\mathcal{L}f^{(k)}\right]\!\left(t^{(m)},\,x^{(m)}\right).$ (16)
The factor $T$ accounts for the uniform measure over $[0,T]$.
Residual and loss.

### 20 Mar 2026 v2

@@ Abstract
We extend the Weak Adversarial Neural Pushforward Method (WANPM) to fractional Fokker-Planck equations, in which the classical Laplacian diffusion operator is replaced by the fractional Laplacian of order $\alpha\in(0,2]$. The solution distribution is represented as the pushforward of a simple base distribution through a neural network, and the weak formulation is discretized entirely via Monte Carlo sampling without any temporal mesh. A key computational advantage is that plane-wave test functions are eigenfunctions of the fractional Laplacian, making the operator cost identical to that of classical diffusion for any $\alpha$. We validate the method on seven benchmark problems with $\alpha=1.5$, spanning one and two spatial dimensions: the steady-state fractional Ornstein–Uhlenbeck (OU) process, a harmonic confining potential, a double-well potential, and a triple-well potential in one dimension, a steady-state 2D double-peak distribution, a time-dependent 2D ring distribution with rotational drift, and a five-dimensional harmonic potential. Each case is benchmarked against particle simulations using symmetric $\alpha$-stable Lévy increments, and robust statistics confirm close agreement throughout. The method is mesh-free, requires no density evaluation or non-local quadrature, and provides a promising foundation for high-dimensional anomalous diffusion solvers.
@@ 1 Introduction
Solving the fFPE in high dimensions poses formidable challenges. Classical finite-element or finite-difference discretisations of the fractional Laplacian require dense stiffness matrices whose assembly and storage become prohibitive even in modest dimensions. Spectral methods are effective in simple geometries but do not readily extend to curved or unbounded domains. Deep learning approaches, which have recently achieved remarkable success for classical high-dimensional PDEs, must contend with two additional obstacles in the fractional setting: (i) pointwise neural-network representations of the solution density require computing the fractional Laplacian of the network output, which involves non-local integral operators and is expensive to evaluate; and (ii) the solution distribution may have heavy algebraic tails that a neural network density must resolve over an exponentially large effective domain.
In a companion paper [3] we introduced the Weak Adversarial Neural Pushforward Method (WANPM) for classical Fokker-Planck equations. The core idea is to represent the solution distribution as the pushforward of a tractable base distribution through a neural map $F_{\bm{\vartheta}}$, and to train $F_{\bm{\vartheta}}$ adversarially against a family of plane-wave test functions so that the induced distribution satisfies the weak form of the FPE. This approach avoids explicit density evaluation, density normalisation, and second-order backpropagation through the solution network. The pushforward map $F_{\bm{\vartheta}}$ can take a base distribution of arbitrary dimension $d$ as input and map it to the $n$-dimensional target space, providing representational flexibility that is particularly valuable for complex or heavy-tailed target distributions.
@@
    We derive the weak formulations of both the time-dependent and steady-state fFPE and show how they can be evaluated as Monte Carlo expectations with respect to the pushforward distribution, without discretising time.
@@
    We establish the theoretical steady-state distribution for the fractional Ornstein–Uhlenbeck process and its connection to $\alpha$-stable laws, and describe the fractional Laplacian eigenfunction property that makes plane-wave test functions exact and computationally efficient for the fractional operator.
@@
    We implement and validate the method on seven benchmark problems with $\alpha=1.5$: the fractional Ornstein–Uhlenbeck steady state, a harmonic confining potential, a double-well potential, and a triple-well potential in one dimension; a steady-state 2D double-peak distribution; a time-dependent 2D ring distribution with rotational drift; and a five-dimensional harmonic potential.
@@
The remainder of the paper is organized as follows. Section 2 presents the fractional Fokker-Planck equation, the steady-state theory, and the weak formulations for both the transient and steady-state problems. Section 3 describes the neural pushforward architecture and the adversarial training procedure. Section 4 presents the particle simulation used as a ground-truth benchmark. Section 5 reports the seven numerical experiments. Section 6 discusses limitations and future directions.
@@ 2.1 Strong Form
For the numerical experiments of this paper we always take $\alpha=1.5$, so that the fractional Laplacian $(-\Delta)^{3/4}$ lies strictly between the first-order and the classical second-order differential operators in terms of its smoothing properties. The solution develops algebraically decaying tails characteristic of $\alpha$-stable distributions, in contrast to the Gaussian tails arising in the classical ($\alpha=2$) case.
Remark 1 (Normalisation convention).
Our convention sets the coefficient of $(-\Delta)^{\alpha/2}$ to unity, so the characteristic function of the Lévy increment over a unit time interval is $e^{-|\xi|^{\alpha}}$. The fractional diffusion coefficient can be absorbed into the wavenumber normalisation of the test functions without loss of generality.
2.2 Steady-State Theory
When the drift is the negative gradient of a confining potential, $b(x)=-V^{\prime}(x)$, and the process is ergodic, equation (2) admits a unique steady-state distribution $\rho_{\infty}$ satisfying
$(-\Delta_{x})^{\alpha/2}\rho_{\infty}+\frac{d}{dx}[b(x)\rho_{\infty}]=0.$ (3)
We now establish the steady-state distribution analytically for the linear (Ornstein–Uhlenbeck) case, and explain why no analogous closed form exists for nonlinear drifts.
Fractional Ornstein–Uhlenbeck steady state.
For the linear drift $b(x)=-\theta(x-\mu)$ with $\theta>0$, the SDE (1) is the fractional Ornstein–Uhlenbeck (fOU) process. Taking the Fourier transform of (3) and using $\widehat{(-\Delta)^{\alpha/2}u}=|\xi|^{\alpha}\hat{u}$ together with $\widehat{\partial_{x}[(x-\mu)\rho_{\infty}]}(\xi)=-i\partial_{\xi}[\xi\hat{\rho}_{\infty}]+i\mu\xi\hat{\rho}_{\infty}$ yields (after shifting to $\mu=0$ for clarity)
$|\xi|^{\alpha}\hat{\rho}_{\infty}(\xi)=-\theta\frac{\partial}{\partial\xi}\bigl[\xi\hat{\rho}_{\infty}(\xi)\bigr].$ (4)
This is a first-order linear ODE in $\xi$. Substituting the $\alpha$-stable ansatz $\hat{\rho}_{\infty}(\xi)=e^{-c^{\alpha}|\xi|^{\alpha}}$ into (4) gives
$c^{\alpha}|\xi|^{\alpha}e^{-c^{\alpha}|\xi|^{\alpha}}=\theta\bigl(1+\alpha c^{\alpha}|\xi|^{\alpha}\bigr)e^{-c^{\alpha}|\xi|^{\alpha}},$
which is satisfied for all $\xi$ if and only if $c^{\alpha}=1/(2\theta)$, i.e. $c=(2\theta)^{-1/\alpha}$. Hence
$\log\hat{\rho}_{\infty}(\xi)=-\frac{|\xi|^{\alpha}}{2\theta}.$ (5)
Proposition 1 (Fractional OU steady state).
For the fOU process with drift $b(x)=-\theta(x-\mu)$, $\theta>0$, and stability index $\alpha\in(0,2)$, the unique steady-state distribution is the symmetric $\alpha$-stable law $S_{\alpha}\!\left((2\theta)^{-1/\alpha},0,\mu\right)$, whose characteristic function is
$\hat{\rho}_{\infty}(\xi)=\exp\!\left(-\frac{|\xi|^{\alpha}}{2\theta}\right)e^{i\mu\xi}.$
For $\alpha=2$ this reduces to the Gaussian $\mathcal{N}(\mu,1/(2\theta))$.
The density of $S_{\alpha}((2\theta)^{-1/\alpha},0,\mu)$ has no elementary closed form for general $\alpha\in(0,2)$, but its tail behavior is known analytically [5]:
$\rho_{\infty}(x)\sim\frac{C_{\alpha}}{(2\theta)\,|x-\mu|^{1+\alpha}}\qquad\text{as }|x-\mu|\to\infty,$ (6)
for an explicit constant $C_{\alpha}>0$. In particular, the distribution has infinite variance for all $\alpha<2$, which is the hallmark of anomalous diffusion and motivates our use of robust statistics rather than second-moment comparisons in Section 5.
General gradient drift.
For a non-linear gradient drift $b(x)=-V^{\prime}(x)$, no closed-form steady state is available for $\alpha<2$. In the classical case ($\alpha=2$, standard Brownian driving noise with diffusivity $\sigma^{2}/2$), the steady state is the Boltzmann distribution $\rho_{\infty}(x)\propto\exp(-V(x)/(\sigma^{2}/2))$. This explicit formula arises because the detailed-balance condition can be integrated exactly to cancel the drift and diffusion contributions. For $\alpha<2$, detailed balance takes the Fourier-space form
$|\xi|^{\alpha}\hat{\rho}_{\infty}(\xi)=i\xi\,\widehat{b\cdot\rho_{\infty}}(\xi),$ (7)
which is a non-local integral equation coupling all frequency modes through the convolution $\widehat{b\cdot\rho_{\infty}}$. No general closed-form inversion is known for non-linear $b$, which is precisely why neural methods benchmarked against long-time particle simulations are valuable for studying fractional steady-state distributions.
2.3 Weak Formulation of the Transient Problem
Let $f(t,x)$ be a smooth test function decaying sufficiently rapidly as $|x|\to\infty$. Multiplying (2) by $f$ and integrating by parts in both space and time over $\mathbb{R}\times[0,T]$ yields
@@ 2.3 Weak Formulation of the Transient Problem
The fractional term $(-\Delta_{x})^{\alpha/2}f$ is well-defined for any smooth $f$ via the Fourier identity $|\xi|^{\alpha}\hat{f}$, and the integration-by-parts step is valid whenever $\rho$ has sufficient integrability.
@@
Equation (8) decomposes into three Monte Carlo-evaluable expectations: a terminal expectation at $t=T$, an initial-condition expectation at $t=0$ (which can be computed from samples of $\rho_{0}$ independently of the learned network), and an interior expectation integrated uniformly over $[0,T]$. This three-term structure is identical to that of the classical FPE [3] up to the replacement of $\partial^{2}/\partial x^{2}$ by $(-\Delta_{x})^{\alpha/2}$ in the operator $\mathcal{L}$. In particular, the fractional and classical problems have the same computational graph; only the scalar multiplier $|\mathbf{w}|^{\alpha}$ in the loss changes with $\alpha$.
2.4 Weak Formulation of the Steady-State Problem
For the steady-state problem (3), multiplying by a smooth test function $f(x)$ and integrating by parts yields the stationarity condition
$\mathbb{E}_{\rho_{\infty}}\!\left[\mathcal{L}_{\mathrm{ss}}f\right]=0\qquad\text{for all smooth }f,$ (10)
where the adjoint operator on the test function is
$\mathcal{L}_{\mathrm{ss}}f(x)=-(-\Delta_{x})^{\alpha/2}f(x)+b(x)\frac{df}{dx}.$ (11)
For sinusoidal test functions $f(x)=\sin(wx+b)$, this becomes
$\mathcal{L}_{\mathrm{ss}}f(x)=-|w|^{\alpha}\sin(wx+b)+b(x)\,w\cos(wx+b),$ (12)
where the fractional term reduces to a simple scalar multiplication via the eigenfunction property. The adversarial training objective is then
$\min_{{\bm{\vartheta}}}\max_{\{w^{(k)},\,b^{(k)}\}}\frac{1}{K}\sum_{k=1}^{K}\Bigl(\mathbb{E}_{\rho_{\bm{\vartheta}}}\bigl[\mathcal{L}_{\mathrm{ss}}f^{(k)}\bigr]\Bigr)^{2},$ (13)
where $\rho_{\bm{\vartheta}}=G_{\bm{\vartheta}}{}_{\#}\pi_{\mathrm{base}}$ is the distribution of the pushforward samples.
@@ 2.5 Eigenfunction Property of Plane-Wave Test Functions
A spatial plane wave $\phi_{\mathbf{w}}(x)=\sin(\mathbf{w}\cdot x+\varphi)$ satisfies
@@
The practical consequence is that the fractional part of $\mathcal{L}f^{(k)}$ costs nothing extra to evaluate compared with classical diffusion: it is simply a multiplication by the scalar $|\mathbf{w}^{(k)}|^{\alpha}$. The entire fFPE loss computation therefore has the same computational complexity as the classical FPE loss, regardless of $\alpha$. This is the key property that makes the plane-wave test function choice especially attractive for fractional operators.
@@ 3 Weak Adversarial Neural Pushforward Method
3.1 Pushforward Network Architecture for the Transient Problem
@@
$F_{\bm{\vartheta}}:[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{d}\to\mathbb{R}^{n},$ (15)
@@
where $\tilde{F}_{\bm{\vartheta}}:\mathbb{R}^{1+n+d}\to\mathbb{R}^{n}$ is a fully connected feedforward network with three hidden layers of width 128 and Tanh activation functions. At $t=0$, the map reduces to the identity on $x_{0}$, so $F_{\bm{\vartheta}}(0,x_{0},\mathbf{r})=x_{0}\sim\rho_{0}$ for any values of the network parameters—the initial condition is satisfied exactly by construction, with no penalisation term required in the loss. The $\sqrt{t}$ prefactor is appropriate for diffusion-driven processes: the typical displacement of an $\alpha$-stable Lévy process over time $t$ scales as $t^{1/\alpha}$, and for $\alpha$ near 2 (and exactly for $\alpha=2$) this is well-approximated by $\sqrt{t}$ at short times.
The base distribution dimension $d$ can be chosen larger than the target dimension $n$, providing additional representational capacity without requiring Jacobian-determinant computations. In our experiments we use $d=5$ for the harmonic and steady-state cases, and $d=8$ for the double-well and triple-well cases where the target distribution is multimodal.
3.2 Pushforward Network Architecture for the Steady-State Problem
For the steady-state problem, the pushforward network has no time input and reduces to a map $G_{\bm{\vartheta}}:\mathbb{R}^{d}\to\mathbb{R}^{n}$, parametrized simply as
$G_{\bm{\vartheta}}(\mathbf{r})=\tilde{G}_{\bm{\vartheta}}(\mathbf{r}),$ (17)
where $\tilde{G}_{\bm{\vartheta}}$ is a fully connected feedforward network with three hidden layers of width 128 and Tanh activations, and $\mathbf{r}\sim\mathcal{N}(0,I_{d})$. Because there is no initial condition to enforce, no $\sqrt{t}$ prefactor is needed, and the network is free to map the Gaussian base distribution to any target shape, including heavy-tailed $\alpha$-stable distributions.
Remark 3 (Overparameterised base distributions).
Allowing the base dimension $d>n$ precludes the use of invertible (normalizing flow) architectures, which require $d=n$ for the change-of-variables formula to hold. This is not a limitation but a feature: the extra latent dimensions provide additional representational capacity without requiring Jacobian-determinant computations. For heavy-tailed targets such as $\alpha$-stable distributions, invertible architectures with $d=n$ may be insufficient, while the overparameterised pushforward can allocate multiple latent directions to represent different features of the tail.
@@ 3.3 Plane-Wave Test Functions
where the last identity uses the eigenfunction property (14). No automatic differentiation through the test function is required. The operator (9) evaluated on the test function is therefore
@@
For the steady-state case, the same formula holds with $\kappa^{(k)}=0$ and the time variable dropped.
3.4 Monte Carlo Discretisation of the Weak Form
The three terms in the weak residual (8) are each approximated by Monte Carlo estimates. We sample $M_{T}$ terminal tuples $(x_{0},\mathbf{r})$ with $x_{0}\sim\rho_{0}$ and $\mathbf{r}\sim\mathcal{N}(0,I_{d})$ to form $\hat{E}_{T}^{(k)}$; $M_{0}$ initial samples $\tilde{x}^{(i)}\sim\rho_{0}$ to form $\hat{E}_{0}^{(k)}$; and $M$ interior triples $(t^{(m)},x_{0}^{(m)},\mathbf{r}^{(m)})$ with $t^{(m)}\sim\mathcal{U}[\epsilon,T]$, $x_{0}^{(m)}\sim\rho_{0}$, $\mathbf{r}^{(m)}\sim\mathcal{N}(0,I_{d})$ to form $\hat{E}^{(k)}$. The factor $T$ in the interior estimate accounts for the uniform measure over $[0,T]$.